# Learning telic-controllable state representations

### v1

@@ Abstract
Computational accounts of purposeful behavior consist of descriptive and normative aspects. The former enable agents to ascertain the current (or future) state of affairs in the world and the latter to evaluate the desirability, or lack thereof, of these states with respect to the agent’s goals. In Reinforcement Learning, the normative aspect (reward and value functions) is assumed to depend on a pre-defined and fixed descriptive one (state representation). Alternatively, these two aspects may emerge interdependently: goals can be, and indeed often are, expressed in terms of state representation features, but they may also serve to shape state representations themselves. Here, we illustrate a novel theoretical framing of state representation learning in bounded agents, coupling descriptive and normative aspects via the notion of goal-directed, or telic, states. We define a new controllability property of telic state representations to characterize the tradeoff between their granularity and the policy complexity capacity required to reach all telic states. We propose an algorithm for learning controllable state representations and demonstrate it using a simple navigation task with changing goals. Our framework highlights the crucial role of deliberate ignorance – knowing what to ignore – for learning state representations that are both goal-flexible and simple. More broadly, our work provides a concrete step towards a unified theoretical view of natural and artificial learning through the lens of goals.
Figure 1: The granularity-complexity tradeoff: within the framework developed in this paper, state representations partition all experiences (blue ellipses) into preference-ordered classes called “telic states” (curved regions), each consisting of all experiences that are (roughly) equivalent with respect to the agent’s goal. Left: an agent with a coarse-grained goal is unable to reach a desired telic state, $S$, which is too distant, in a statistical sense, from its default policy $\pi_{0}$. Right: updating the goal results in a finer-grained, and therefore more controllable, telic state representation, where all states can be reached via simple policy updates.
@@ 1 Introduction
The role of goals in shaping how agents learn to represent their environment is becoming the focus of increased attention in both cognitive science (Molinaro & Collins, 2023; Muhle-Karbe et al., 2023; Radulescu et al., 2019) and reinforcement learning (Eysenbach et al., 2022; Florensa et al., 2018; Wang et al., 2024). A fundamental open problem however, is how to learn useful task representations when goals are unstable and computational resources are bounded. For example, consider a rodent navigating a complex maze with changing reward contingencies (Krausz et al., 2023), or a robot trained to do various object manipulation tasks using only binary and sparse rewards (Andrychowicz et al., 2017). How can such learning agents represent their environments in ways that facilitate adaptation to shifting goals using limited computational resources? While several heuristic methods have been proposed to address this problem, here we describe a principled approach, leveraging a recently proposed theoretical framework of goal-directed, or telic, state representation learning (Amir et al., 2023). Drawing on control and information theoretic principles, we define novel property, called telic-controllability, characterizing the ability of a complexity bounded agent to reach all states within a given telic state representation. We propose a state representation learning algorithm and demonstrate it by showing how complexity bounded agents can learn a telic-controllable state representation for a simple navigation task with a shifting goal.
@@ 2.1 Telic states as goal-equivalent experiences
We assume the setting of a perception-action cycle, i.e., sequences of observation-action pairs representing the flow of information between agent and environment. We denote by $\mathcal{O}$ and $\mathcal{A}$ the set of possible observations and actions, respectively. An experience sequence, or experience for short, is a finite sequence of observation-action pairs: $h=o_{1},a_{1},o_{2},a_{2},...,o_{n},a_{n}$. For every non-negative integer, $n\geq 0$, we denote by $\mathcal{H}_{n}\equiv(\mathcal{O}\times\mathcal{A})^{n}$ the set of all experiences of length $n$. The collection of all finite experiences is denoted by $\mathcal{H}=\cup_{n=1}^{\infty}\mathcal{H}_{n}$. In non-deterministic settings, it will be useful to consider distributions over experiences rather than individual experiences themselves and we denote the set of all probability distributions over finite experiences by $\Delta(\mathcal{H})$. Following Bowling et al. (2022), we define a goal as a binary preference relation over experience distributions. For any pair of experience distributions, $A,B\in\Delta(\mathcal{H})$, we write $A\succeq_{g}B$ to indicate that experience distribution $A$ is weakly preferred by the agent over $B$, i.e., that $A$ is at least as desirable as $B$, with respect to goal $g$. When $A\succeq_{g}B$ and $B\succeq_{g}A$ both hold, $A$ and $B$ are equally preferred with respect to $g$, denoted as $A\sim_{g}B$. We observe that $\sim_{g}$ is an equivalence relation, i.e., it satisfies the following three properties, for any $A,B,C\in\Delta(\mathcal{H})$: (1) $A\sim_{g}A$ (reflexivity); (2) $A\sim_{g}B\implies B\sim_{g}A$ (symmetry); and (3) $(A\sim_{g}B)\wedge(B\sim_{g}C)\implies A\sim_{g}C$ (transitivity). Therefore, every goal induces a partition of $\Delta(\mathcal{H})$ into disjoint sets of equally desirable experience distributions. For goal $g$, we define the goal-directed, or telic, state representation, $\mathcal{S}_{g}$, as the partition of experience distributions into equivalence classes it induces:
$\mathcal{S}_{g}=\Delta(\mathcal{H})/\sim_{g}.$ (1)
In other words, each telic state represents a generalization over all equally desirable experience distributions. This definition captures the intuition that agents need not distinguish between experiences that are equivalent, in a statistical sense, with respect to their goal. Furthermore, since different telic states are, by definition, non-equivalent with respect to $\succeq_{g}$, the goal $g$ also determines whether a transition between any two telic states brings the agent in closer alignment to, or further away from its goal. See Amir et al. (2023) for additional details.
@@ 2.2 Experience features and discrimination sensitivity
A natural way of representing goals, i.e., preferences over experience distributions, is by comparing the likelihood that experiences generated from different distributions will belong to some subset $\Phi_{g}\subset\mathcal{H}$ representing some desired property of experiences. In other words, for two experience distributions, $A$ and $B$, the agent will prefer the one that is more likely to generate experiences belonging to $\Phi_{g}$:
$A\succeq_{g}B:\sum_{h\in\Phi_{g}}A(h)\geq\sum_{h\in\Phi_{g}}B(h).$ (2)
For example, for a goal of solving a maze, $\Phi_{g}$ might be the set of all experiences, i.e., path trajectories, that reach the exit. Thus, experience distribution $A$ would be preferred over $B$ if it is more likely to generate trajectories that reach the exit. Importantly however, Eq. 2 implies that $A$ and $B$ are equivalent only when $\sum_{h\in\Phi_{g}}A(h)$ and $\sum_{h\in\Phi_{g}}B(h)$ are precisely equal, which is unlikely in realistic, noisy environments. A more reasonable assumption is that agents can discriminate sampling likelihoods at some finite sensitivity level, $\epsilon>0$, such that:
@@
In the maze example, this means that two trajectory distributions are considered equivalent if their likelihood of generating exit-reaching trajectories is within an $\epsilon-$neighbourhood of each others. As we shall see in the following sections, the discrimination sensitivity parameter, $\epsilon$, determines the granularity of the telic state representation, which in turn determines the complexity of policies needed to reach different telic states.
@@ 2.3 Telic-controllability
As detailed in appendix. A, this mapping between polices and telic states provides a unified account of goal-directed learning in terms the statistical distance between policy-induced distributions and desired telic states. To explore this notion, we introduce a new property – telic-controllability – that plays a central role in the following sections. A representation is called telic-controllable if any state can be reached using a finite number, $N$, of complexity-limited policy updates, starting from the agent’s default policy, $\pi_{0}$, where the complexity of a policy update step is quantified by the Kullback-Leibler (KL) divergence between the post and pre-update step policies. Formally, we define:
@@ 2.4 State representation learning algorithm
A central feature of our approach is the coupling, or duality, it establishes between goals and state representations. In this section, we leverage this duality to propose an algorithm for learning a telic-controllable state representation, or, equivalently, forming a new goal that produces such a state representation. The algorithm receives as inputs the agent’s current goal, $g$ (represented, e.g., by an ordered set of desired experience features), and default policy, $\pi_{0}$, along with its policy complexity capacity, $\delta$, and the discrimination sensitivity parameter $\epsilon$. It output consists of a new goal $g^{\prime}$ such that $\mathcal{S}_{g^{\prime}}$ is telic controllable with respect to $\pi_{0}$ and $\delta$. The main idea is to split any unreachable telic states, i.e., those that cannot be reached from the agent’s default policy with updates that satisfy its complexity constraints. State splitting is accomplished by generating a new intermediate state lying between the agent’s default policy induced distribution, $P_{\pi_{0}}$ and the distribution closest to it, in the KL sense, in the unreachable state, $S$, denoted $P^{*}$. The intermediate state, $S_{M}$, is defined as the set of all distributions that are $\epsilon$-equivalent to $P_{M}$ (Eq. 3), where $P_{M}$ is the convex combination of $P^{*}$ and $P_{\pi_{0}}$ lying at a KL distance of $\delta$ from $P_{\pi_{0}}$. After generating the new state, $S_{M}$, the goal is updated to reflect the proper order between the default policy state $S_{0}$, the intermediate state $S_{M}$ and the originally unreachable state $S$, such that elements of $S_{M}$ are between $S_{0}$ and $S$ in terms of preference. Pseudocode for the learning algorithm is provided in appendix. B.
@@ 3 Illustrative example: dual goal navigation task
In this section, we illustrate the proposed state representation framework and learning algorithm using a simple navigation task in which an agent performs a one dimensional random walk, starting at location $x_{0}=0$, and aiming to reach one of two non overlapping regions of interest after a fixed number, $T=30$, of steps. The agent’s policy is defined as a stochastic mapping between its current and next position and is parameterized by the mean and standard deviation ($\mu$ and $\sigma$, respectively) of a Gaussian update step: $\pi(x_{t+1}|x_{t};\mu,\sigma)=x_{t}+\eta_{t},\quad\eta_{t}\sim\mathcal{N}(\mu,\sigma).$ For brevity, we denote by $\pi(\mu,\sigma)$ a policy with a $\mathcal{N}(\mu,\sigma)$ distributed noise term. See appendix C for a graphical illustration of the task and sample trajectories for different policies.
@@
To account for goal-directed behavior, we define a right and a left region of interest, $R$ and $L$, consisting of unit radius segments centered around $x_{R}=2$ and $x_{L}=-2$ respectively. Thus, $R=[R_{1},R_{2}]=[1,3]$ and $L=[L_{1},L_{2}]=[-3,-1]$. For the purpose of this example, we assume that the agent wants to reach $R$, but avoid $L$, at time $T$. For example, for a rodent navigating a narrow corridor, $R$ and $L$ may indicate segments of the corridor where a reward (e.g., food) and a punishment (e.g., air puff) are administered, respectively. We can express the agent’s goal in terms of preferences over policies by defining $\Delta P(\mu,\sigma)=p(x_{T}\in R\mid\mu,\sigma)-p(x_{T}\in L\mid\mu,\sigma)$ as the difference between the probabilities that the agent will reach regions $R$ and $L$ at time $T$, with a policy $\pi(\mu,\sigma)$. The agent’s goal can now be defined as a preference for policies with higher $\Delta P$ values. However, as explained in section 2.2 above, due to the agent’s finite discrimination resolution, it can only detect whether $\Delta P$ is above or below the resolution threshold, $\epsilon$. Thus, using Eq. 4, the agent’s goal, $g$, can be expressed as the following preference relation over policies:
@@
where first term on the r.h.s. of Eq. 7 captures the desirability of $R$ – the agent prefers policies that have a probability higher than $\epsilon$ of reaching $R$ over ones that do not; while the second term captures the undesirability of $L$ – the agent prefers policies that have a probability lower than $\epsilon$ to reach $L$ than ones that do not. The discrimination threshold, $\epsilon$, thus determines the borders between the resulting telic states. The telic state representation for the goal $g$ defined by Eq. 7, and a threshold parameter of $\epsilon=0.1$ is visualized in Fig. 2 (top left). Telic state $S_{R}$ ($S_{L}$), is shown as a colored region bounded by a dotted green (red) line, consisting of all policies that are more (less) likely to reach $R$ than $L$ by a probability margin of $\epsilon$ or more. Policies that are roughly equally likely to reach $R$ or $L$, i.e., whose difference in $\Delta P$ is smaller than $\epsilon$, constitute an additional “default” telic state, $S_{0}$ (shown as a teal background), in which the agent is agnostic as to which region is it more likely to reach.
@@
$S_{R}$ =$=\left\{(\mu,\sigma)\bigg{|}\frac{1}{2}\left(\operatorname{erf}{\frac{R_{1}-T\mu}{{\sqrt{2T}}\sigma}}-\operatorname{erf}{\frac{R_{2}-T\mu}{{\sqrt{2T}}\sigma}}\right)-\frac{1}{2}\left(\operatorname{erf}{\frac{L_{1}-T\mu}{{\sqrt{2T}}\sigma}}-\operatorname{erf}{\frac{L_{2}-T\mu}{{\sqrt{2T}}\sigma}}\right)\geq\epsilon\right\},$ (9)
@@
with similar expressions for $S_{L}$ and $S_{0}$. To illustrate the notion of telic-controllability (Eq. 5) using this representation, we define the complexity of any policy $\pi(\mu,\sigma)$ with respect to the agent’s default policy, $\pi_{0}(\mu_{0},\sigma_{0})$, as the KL divergence, per time step, between them:
$D_{KL}(\pi\|\pi_{0})=\log\frac{\sigma_{0}}{\sigma}+\frac{\sigma^{2}+(\mu-\mu_{0})^{2}}{2\sigma_{0}^{2}}-\frac{1}{2},$ (10)
@@
which is the KL divergence between the two Gaussians $\mathcal{N}(\mu,\sigma)$ and $\mathcal{N}(\mu_{0},\sigma_{0})$. The contour lines in the first three panels of Fig. 2 (top & bottom left) show isometric policy complexity levels for an agent with a capacity of $\delta=1$ bit per time step, and a default policy $\pi_{0}(\mu_{0}=0,\sigma_{0}=1)$. Initially, both telic states, $S_{R}$ and $S_{L}$, lie within the range of the agent’s policy complexity capcity (top left). The policies in $S_{R}$ and $S_{L}$ that are closest in the $KL$ sense to $\pi_{0}$ (green and red dots, respectively), both lie within a range of less than $\delta$ from $\pi_{0}$, i.e., the state representation is telic-controllable. When the center of $R$ shifts from $x_{R}=2$ to $x_{R}=2.5$ (top right), telic state $S_{R}$ is no longer within complexity range $\delta$ from $\pi_{0}$ and the state representation becomes non-controllable. To address this (bottom left), the state representation learning algorithm described in 2.4, splits $S_{R}$ by adding an intermediate telic state $S_{M}$ (orange), centered around the policy closest to $S_{R}$ that is still within a KL-range of $\delta$ from $\pi_{0}$ (yellow dot). This changes the shape of $S_{R}$ and $S_{L}$ since now the probability of reaching each of the three telic states, $S_{R},S_{L}$ and $S_{M}$, is defined in with respect to the two others, e.g., $S_{M}=\left\{(\mu,\sigma)|\Delta P_{M}(\mu,\sigma)\geq\epsilon\right\}$ where $\Delta P_{M}=p(x_{T}\in M\mid\mu,\sigma)-\max\{p(x_{T}\in L\mid\mu,\sigma),p(x_{T}\in R\mid\mu,\sigma)\}$, and similarly for $S_{R}$ and $S_{L}$. Since $\pi_{M}$ is, by construction, within a KL range of $\delta$ from $\pi_{0}$, the agent can reach $S_{M}$ by updating its default policy to $\pi_{M}$ (bottom right), bringing $S_{R}$ into reach again. Hence, the new state representation, consisting of $S_{0},S_{L},S_{M}$ and $S_{R}$, is tellic-controllable. Finally, Fig. 3 illustrates the granularity-complexity tradeoff: the granularity of the state representation, quantified as $-\log(\epsilon)$ (abscissa), controls the complexity capacity required to reach each state (ordinate). Finer-grained representations are generally more controllable. For a granularity level of $\epsilon=0.1$ (gray vertical line), only $S_{L}$ and $S_{M}$ are reachable from $\pi_{0}(0,1)$ under a complexity capacity of $\delta=1$ (gray horizontal line).

### v2

@@ Abstract
Computational descriptions of purposeful behavior comprise both descriptive and normative aspects. The former are used to ascertain current (or future) states of the world and the latter to evaluate the desirability, or lack thereof, of these states under some goal. In Reinforcement Learning, the normative aspect (reward and value functions) is assumed to depend on a predefined and fixed descriptive one (state representation). Alternatively, these two aspects may emerge interdependently: goals can be, and indeed often are, approximated by state-dependent reward functions, but they may also shape the acquired state representations themselves. Here, we present a novel computational framework for state representation learning in bounded agents, where descriptive and normative aspects are coupled through the notion of goal-directed, or telic, states. We introduce the concept of telic controllability to characterize the tradeoff between the granularity of a telic state representation and the policy complexity required to reach all telic states. We propose an algorithm for learning controllable state representations, illustrating it using a simple navigation task with shifting goals. Our framework highlights the crucial role of deliberate ignorance – knowing which features of experience to ignore – for learning state representations that balance goal flexibility and policy complexity. More broadly, our work advances a unified theoretical perspective on goal-directed state representation learning in natural and artificial agents.
Figure 1: The granularity-complexity tradeoff: within the framework developed in this paper, state representations partition all experiences (blue ellipses) into preference-ordered classes called “telic states” (curved regions), each consisting of all experience distributions that are (roughly) equivalent with respect to the agent’s goal. Left: an agent with a coarse-grained goal is unable to reach a desired telic state, $S$, which is too distant, in a statistical sense, from its default policy $\pi_{0}$. Right: by refining the goal the agent can acquire a more controllable state representation, where all telic states can be reached via simple policy update steps.
@@ 1 Introduction
How do goals shape the state representations acquired by learning agents? This question is the focus of growing attention in both cognitive science (Molinaro & Collins, 2023; Muhle-Karbe et al., 2023; Radulescu et al., 2019) and reinforcement learning (Eysenbach et al., 2022; Florensa et al., 2018; Wang et al., 2024a). A fundamental open problem however, is how to acquire useful state representations when goals are changing and computational resources are bounded. For example, consider a rodent navigating a complex maze with changing reward contingencies (Krausz et al., 2023), or a robot trained to do various object manipulation tasks with sparse rewards (Andrychowicz et al., 2017). How should such learning agents represent their environments in ways that facilitate adaptation to shifting goals using limited computational resources? Prior works have addressed the problem of efficient state representation learning using bisimulation (Zhang et al., 2020; Wang et al., 2024b) or option-based methods (Abel et al., 2020). While these approaches can provide efficient heuristics for state abstraction in Markovian settings, our focus here is on the fundamental relationship between the granularity of state representations, which may be non-Markovian, and the policy complexity resources needed to efficiently utilize them. We present a principled approach to this problem, leveraging a recently proposed theoretical framework of goal-directed, or telic, state representation learning (Amir et al., 2023). We define a novel property, telic-controllability, characterizing the ability of reaching all states within a given telic state representation using complexity bound policies. We describe a telic state representation learning algorithm and illustrate it using a simple navigation task by showing how complexity bounded agents can learn a telic-controllable state representation that can adapt to shifting goals.
@@ 2.1 Telic states as goal-equivalent experiences
We assume the setting of a perception-action cycle, i.e., sequences of observation-action pairs representing the flow of information between agent and environment. We denote by $\mathcal{O}$ and $\mathcal{A}$ the set of possible observations and actions, respectively. An experience sequence, or experience for short, is a finite sequence of observation-action pairs: $h=o_{1},a_{1},o_{2},a_{2},...,o_{n},a_{n}$. For every non-negative integer, $n\geq 0$, we denote by $\mathcal{H}_{n}\equiv(\mathcal{O}\times\mathcal{A})^{n}$ the set of all experiences of length $n$. The collection of all finite experiences is denoted by $\mathcal{H}=\cup_{n=1}^{\infty}\mathcal{H}_{n}$. In non-deterministic settings, it will be useful to consider distributions over experiences rather than individual experiences themselves and we denote the set of all probability distributions over finite experiences by $\Delta(\mathcal{H})$. Following Bowling et al. (2022), we define a goal as a binary preference relation over experience distributions. For any pair of experience distributions, $A,B\in\Delta(\mathcal{H})$, we write $A\succeq_{g}B$ to indicate that experience distribution $A$ is weakly preferred by the agent over $B$, i.e., that $A$ is at least as desirable as $B$, with respect to goal $g$. When $A\succeq_{g}B$ and $B\succeq_{g}A$ both hold, $A$ and $B$ are equally preferred with respect to $g$, denoted as $A\sim_{g}B$. We observe that $\sim_{g}$ is an equivalence relation, i.e., it satisfies the following three properties, for any $A,B,C\in\Delta(\mathcal{H})$:
  • Reflexivity: $A\sim_{g}A$ for all $A\in\Delta(\mathcal{H}).$
  • Symmetry: $A\sim_{g}B$ implies $B\sim_{g}A$ for all $A,B\in\Delta(\mathcal{H}).$
  • Transitivity: if $A\sim_{g}B$ and $B\sim_{g}C$ then $A\sim_{g}C$ for all $A,B,C\in\Delta(\mathcal{H}).$
Therefore, every goal induces a partition of $\Delta(\mathcal{H})$ into disjoint sets of equally desirable experience distributions. For goal $g$, we define the goal-directed, or telic, state representation, $\mathcal{S}_{g}$, as the partition of experience distributions into equivalence classes it induces: $\mathcal{S}_{g}=\Delta(\mathcal{H})/\sim_{g}$. In other words, each telic state represents a generalization over all equally desirable experience distributions. This definition captures the intuition that agents need not distinguish between experiences that are equivalent, in a statistical sense, with respect to their goal. Furthermore, since different telic states are, by definition, non-equivalent with respect to $\succeq_{g}$, the goal $g$ also determines whether a transition between any two telic states brings the agent in closer alignment to, or further away from its goal. See Amir et al. (2023) for additional details.
@@ 2.2 Experience features and discrimination sensitivity
A natural way of representing goals, i.e., preferences over experience distributions, is by comparing the likelihood that experiences generated from different distributions will belong to some subset $\Phi_{g}\subset\mathcal{H}$ representing some desired property of experiences. For example, for a goal of solving a maze, $\Phi_{g}$ might be the set of all experiences, i.e., path trajectories, that reach the exit. Formally, for two experience distributions, $A$ and $B$, the agent will prefer the one that is more likely to generate experiences belonging to $\Phi_{g}$:
$A\succeq_{g}B:\sum_{h\in\Phi_{g}}A(h)\geq\sum_{h\in\Phi_{g}}B(h).$
In the maze example, experience distribution $A$ would be preferred over $B$ if it is more likely to generate trajectories that reach the exit. Importantly, Eq. 2.2 implies that $A$ and $B$ are equivalent only when $\sum_{h\in\Phi_{g}}A(h)$ and $\sum_{h\in\Phi_{g}}B(h)$ are precisely equal, which is unlikely in realistic, noisy environments. A more reasonable assumption is that agents can discriminate sampling likelihoods at some finite sensitivity level, $\epsilon>0$, such that:
@@
In the maze example, this means that two trajectory distributions are considered equivalent if their respective likelihoods of generating exit-reaching trajectories is within $\epsilon$ of each other. As we shall see in the following sections, the discrimination sensitivity parameter, $\epsilon$, controls the tradeoff between the granularity of a telic state representation and the policy complexity needed to reach all telic states.
@@ 2.3 Telic-controllability
As detailed in appendix. A, this mapping between polices and telic states provides an unified account of goal-directed learning in terms the statistical distance between policy-induced distributions and desired telic states. To explore this notion, we introduce a new property – telic-controllability – that plays a central role in the following sections. A representation is called telic-controllable if any state can be reached using a finite number, $N$, of complexity-limited policy updates, starting from the agent’s default policy, $\pi_{0}$, where the complexity of a policy update step is quantified by the Kullback-Leibler (KL) divergence between the post and pre-update step policies. Formally, we have the following:
@@ 2.4 State representation learning algorithm
A central feature of our approach is the duality it establishes between goals and state representations. In this section, we utilize this duality to develop an algorithm for learning a telic-controllable state representation, or, equivalently, finding a goal that produces such a state representation. The algorithm receives as inputs the agent’s current goal, $g$ (represented, e.g., by an ordered set of desired experience features), and default policy, $\pi_{0}$, along with its policy complexity capacity, $\delta$, and the discrimination sensitivity parameter $\epsilon$. It output consists of a new goal $g^{\prime}$ such that $\mathcal{S}_{g^{\prime}}$ is telic controllable with respect to $\pi_{0}$ and $\delta$. The main idea is to split any unreachable telic state, $S$, i.e., one that cannot be reached from $\pi_{0}$ using policy update steps with complexity less than $\delta$. State splitting is accomplished by generating a new, intermediate, telic state, $S_{M}$, lying between the agent’s default policy induced distribution, $P_{\pi_{0}}$, and its information projection on the unreachable telic state, i.e., the distribution $P^{*}\in S$ that is closest to $P_{\pi_{0}}$, in the KL sense. The intermediate telic state, $S_{M}$, is then defined as the set of all distributions that are $\epsilon$-equivalent to $P_{M}$ (Eq. 1), where $P_{M}$ is the convex combination of $P^{*}$ and $P_{\pi_{0}}$ lying at a KL distance of $\delta$ from $P_{\pi_{0}}$. After generating the new state, $S_{M}$, the goal is updated to reflect the proper ordering between the default policy state $S_{0}$, the intermediate state $S_{M}$ and the originally unreachable state $S$, such that elements of $S_{M}$ are between $S_{0}$ and $S$ in terms of preference. Pseudocode for the learning algorithm is provided in appendix. B.
@@ 3 Illustrative example: dual goal navigation task
In this section, we illustrate the proposed state representation framework and learning algorithm using a simple navigation task in which an agent performs a one dimensional random walk, starting at location $x_{0}=0$, with the goal of reaching one of two non overlapping regions of interest after a fixed number, $T=30$, of steps. The agent’s policy is defined as a stochastic mapping between its current and next position and is parameterized by the mean and standard deviation ($\mu$ and $\sigma$, respectively) of a Gaussian update step: $\pi(x_{t+1}|x_{t};\mu,\sigma)=x_{t}+\eta_{t},\quad\eta_{t}\sim\mathcal{N}(\mu,\sigma).$ For brevity, we denote by $\pi(\mu,\sigma)$ a policy with a $\mathcal{N}(\mu,\sigma)$ distributed noise term. A graphical illustration of the task and sample trajectories for different policies is shown in Fig. 2.
Figure 2: Dual goal navigation task: each tile shows 500 one-dimensional random walk trajectories of length $T=30$, generated by a Gaussian policy parameterized by the mean ($\mu$) and standard deviation ($\sigma$) of position update step (x-axis) across time (y-axis). Regions of interest $R$ and $L$ consist of line segments centered around $x_{R}=2$ and $x_{L}=-2$, shown as green and red lines respectively at $T=30$. Trajectories reaching one of the goals are plotted in the corresponding color, illustrating the relationship between policy parameters and goal reaching likelihoods. The default policy, $(\mu_{0},\sigma_{0})=(0,1)$, shown in the center gray tile, is equally likely to reach $R$ and $L$.
@@
To account for goal-directed behavior, we define a right and a left region of interest, $R$ and $L$, consisting of unit radius segments centered around $x_{R}=2$ and $x_{L}=-2$ respectively. Thus, $R=[R_{1},R_{2}]=[1,3]$ and $L=[L_{1},L_{2}]=[-3,-1]$. For the purpose of this example, we assume that the agent wants to reach $R$, but avoid $L$, at time $T$. For example, for a rodent navigating a narrow corridor, $R$ and $L$ may indicate segments of the corridor where a reward (e.g., food) and a punishment (e.g., air puff) are administered, respectively. We can express the agent’s goal in terms of preferences over policies by defining $\Delta P(\mu,\sigma)=p(x_{T}\in R\mid\mu,\sigma)-p(x_{T}\in L\mid\mu,\sigma)$ as the difference between the probabilities that the agent will reach regions $R$ and $L$ at time $T$, with a policy $\pi(\mu,\sigma)$. The agent’s goal can now be defined as a preference for policies with higher $\Delta P$ values. However, as explained in section 2.2 above, due to the agent’s finite discrimination resolution, it can only detect whether $\Delta P$ is above or below the resolution threshold, $\epsilon$. Thus, using Eq. 2, the agent’s goal, $g$, can be expressed by the following preference relation over policies:
@@
where first term on the r.h.s. of Eq. 5 captures the desirability of $R$ – the agent prefers policies that have a probability higher than $\epsilon$ of reaching $R$ over ones that do not; while the second term captures the undesirability of $L$ – the agent prefers policies that have a probability lower than $\epsilon$ to reach $L$ than ones that do not. The discrimination threshold, $\epsilon$, thus determines the borders between the resulting telic states. The telic state representation for the goal $g$ defined by Eq. 5, and a threshold parameter of $\epsilon=0.1$ is visualized in Fig. 3 (top left). Telic state $S_{R}$ ($S_{L}$), is shown as a colored region bounded by a dotted green (red) line, consisting of all policies that are more (less) likely to reach $R$ than $L$ by a probability margin of $\epsilon$ or more. Policies that are roughly equally likely to reach $R$ or $L$, i.e., whose difference in $\Delta P$ is smaller than $\epsilon$, constitute an additional “default” telic state, $S_{0}$ (teal background), in which the agent is agnostic to which region is it more likely to reach.
@@
$S_{R}$ =$=\left\{(\mu,\sigma)\bigg{|}\frac{1}{2}\left(\operatorname{erf}{\frac{R_{1}-T\mu}{{\sqrt{2T}}\sigma}}-\operatorname{erf}{\frac{R_{2}-T\mu}{{\sqrt{2T}}\sigma}}\right)-\frac{1}{2}\left(\operatorname{erf}{\frac{L_{1}-T\mu}{{\sqrt{2T}}\sigma}}-\operatorname{erf}{\frac{L_{2}-T\mu}{{\sqrt{2T}}\sigma}}\right)\geq\epsilon\right\},$
@@
with similar expressions for $S_{L}$ and $S_{0}$. To illustrate the notion of telic-controllability (Eq. 3) using this representation, we define the complexity, $C(\pi)$, of a policy, $\pi(\mu,\sigma)$, with respect to the agent’s default policy, $\pi_{0}(\mu_{0},\sigma_{0})$, as the KL divergence, per time step, between them:
$C(\pi)\equiv D_{KL}(\pi\|\pi_{0}).$
@@
The contour lines in the first three panels of Fig. 3 (top & bottom left) show isometric policy complexity levels for an agent with a complexity capacity of $\delta=1$ bit per time step, and a default policy $\pi_{0}(\mu_{0}=0,\sigma_{0}=1)$. Initially, both telic states, $S_{R}$ and $S_{L}$, lie within the range of the agent’s policy complexity capcity (top left). The policies in $S_{R}$ and $S_{L}$ that are closest in the $KL$ sense to $\pi_{0}$ (green and red dots, respectively), both lie within a range of less than $\delta$ from $\pi_{0}$, i.e., the state representation is telic-controllable. When the center of $R$ shifts from $x_{R}=2$ to $x_{R}=2.5$ (top right), telic state $S_{R}$ is no longer within complexity range $\delta$ from $\pi_{0}$ and the state representation becomes non-controllable. To address this (bottom left), the state representation learning algorithm described in 2.4, splits $S_{R}$ by adding an intermediate telic state $S_{M}$ (orange), centered around the policy closest to $S_{R}$ that is still within a KL-range of $\delta$ from $\pi_{0}$ (yellow dot). This changes the shape of $S_{R}$ and $S_{L}$ since now the probability of reaching each of the three telic states, $S_{R},S_{L}$ and $S_{M}$, is defined in with respect to the two others, e.g., $S_{M}=\left\{(\mu,\sigma)|\Delta P_{M}(\mu,\sigma)\geq\epsilon\right\}$ where $\Delta P_{M}=p(x_{T}\in M\mid\mu,\sigma)-\max\{p(x_{T}\in L\mid\mu,\sigma),p(x_{T}\in R\mid\mu,\sigma)\}$, and similarly for $S_{R}$ and $S_{L}$. Since $\pi_{M}$ is, by construction, within a KL range of $\delta$ from $\pi_{0}$, the agent can reach $S_{M}$ by updating its default policy to $\pi_{M}$ (bottom right), bringing $S_{R}$ into reach again. Hence, the new state representation, consisting of $S_{0},S_{L},S_{M}$ and $S_{R}$, is tellic-controllable. Finally, Fig. 4 illustrates the granularity-complexity tradeoff: the granularity of the state representation, quantified as $-\log(\epsilon)$ (abscissa), controls the complexity capacity required to reach each state (ordinate). Finer-grained representations are generally more controllable. For a granularity level of $\epsilon=0.1$ (gray vertical line), only $S_{L}$ and $S_{M}$ are reachable from $\pi_{0}(0,1)$ under a complexity capacity of $\delta=1$ (gray horizontal line).